OpenStudy (anonymous):

Which of the following polynomials is prime over the set of rational numbers? a.x^2-13x-36 b.x^2-5x-36 c.x^2+15x+36 d.x^2+20x+36

OpenStudy (anonymous):

check to see for each of \[b^2-4ac\] is a perfect square ( 0 counts as a perfect square) if so, rational zeros. if not, irrational zeros so prime over the rationals

OpenStudy (anonymous):

for example in the first one \[13^2-4\times 1\times (-36)=313\] which is not a perfect square.

OpenStudy (anonymous):

where as in the second one \[5^2+4\times 36=169\] which is a perfect square since it is the square of 13